# TextInPlace: Indoor Visual Place Recognition in Repetitive Structures with Scene Text Spotting and Verification

### v1

@@ Abstract
Visual Place Recognition (VPR) is a crucial capability for long-term autonomous robots, enabling them to identify previously visited locations using visual information. However, existing methods remain limited in indoor settings due to the highly repetitive structures inherent in such environments. We observe that scene text typically appears in indoor spaces, serving to distinguish visually similar but different places. This inspires us to propose TextInPlace, a simple yet effective VPR framework that integrates Scene Text Spotting (STS) to mitigate visual perceptual ambiguity in repetitive indoor environments. Specifically, TextInPlace adopts a dual-branch architecture within a local parameter sharing network. The VPR branch employs attention-based aggregation to extract global descriptors for coarse-grained retrieval, while the STS branch utilizes a bridging text spotter to detect and recognize scene text. Finally, the discriminative text is filtered to compute text similarity and re-rank the top-K retrieved images. To bridge the gap between current text-based repetitive indoor scene datasets and the typical scenarios encountered in robot navigation, we establish an indoor VPR benchmark dataset, called Maze-with-Text. Extensive experiments on both custom and public datasets demonstrate that TextInPlace achieves superior performance over existing methods that rely solely on appearance information. The dataset, code, and trained models are publicly available at https://github.com/HqiTao/TextInPlace.
@@ I INTRODUCTION
However, State-of-The-Art (SOTA) VPR methods often struggle in indoor scenarios, suggesting that indoor VPR remains an underexplored research area. Moreover, indoor VPR presents a more convenient and cost-effective alternative compared to traditional indoor localization methods that rely on WiFi or Radio Frequency Identification (RFID). This highlights the importance of further research and innovation to develop robust and efficient VPR techniques tailored specifically for indoor environments. We argue that the suboptimal performance of current VPR methods in indoor environments can be primarily attributed to the repeated structure caused by the repetitive and symmetrical designs of man-made buildings. Unlike outdoor environments, indoor spaces often exhibit high visual similarity across different locations (see Fig. 1), particularly in multi-floor buildings where architectural designs, floor layouts, and decorative elements are often highly repetitive [8] [9].
@@
Inspired by human cognitive patterns, several recent studies [10] [11] [12] [13] proposed leveraging the high-level semantic information embedded in scene text for robot localization. However, these methods also typically rely on off-the-shelf Scene Text Spotting (STS) technology, which frequently becomes a computational bottleneck. Furthermore, they mainly use a predefined dictionary of specific words to filter out discriminative text, which limits their adaptability to diverse environments and requires significant manual effort. Additionally, currently available text-based localization datasets, which are collected with a focus on scene text, remain limited in number. Existing datasets deliberately collect scene text images from close distances and frontal perspectives to facilitate text spotting, but this collection process does not accurately represent the typical scenarios of robot navigation in real-world environments.
@@
To enhance the applicability of VPR in indoor settings, we propose TextInPlace, a unique dual-branch VPR framework that harnesses textual clues for robust spatial verification in highly repetitive indoor environments. Specifically, TextInPlace performs coarse-grained retrieval based on the global descriptor derived from the VPR branch, and executes scene text spotting via a dedicated STS branch applied to both query and retrieved images. Discriminative text is efficiently filtered out based on predefined rules or, more generally, using a large language model (LLM). It then establishes textual semantic correspondence by matching these discriminative scene text, enabling explainable re-ranking of the top-K candidate images initially retrieved by VPR branch. Moreover, the VPR branch and STS branch are combined with a local parameter sharing backbone for optimizing computational efficiency. To address the limitations of existing datasets, we use a panoramic camera to establish a indoor VPR benchmark dataset, called Maze-with-Text. The panoramic images are then cropped to generate VPR data. Each image in the dataset not only contains scene text, but also captures the structural details of the environment.
@@
  • We introduce an LLM-based text filter that flexibly and generically identifies discriminative text from complex scene text.
  • We establish Maze-with-Text, an indoor VPR benchmark dataset to highlight the challenge of repetitive structures in multi-floor buildings, with scene text appearing at various distances and perspectives, thus more closely reflecting robot navigation scenarios.
@@ II-A Visual Place Recognition Network
The VPR network consists mainly of backbone architecture and aggregation layer. Both CNN-based [3] [4] [14] and Transformer-based [15] [16] feature extraction backbones have been widely employed in VPR research, demonstrating their effectiveness. Recently, researchers adopted the vision foundation model DINOv2 [17] as a backbone, exploring both train-free [18] and fine-tuning [19] [20] approaches, and achieving remarkable capabilities. However, vision foundation models are computationally intensive, and relying on large backbones may lead researchers to overlook innovations in other aspects of the VPR problem [21]. The role of the aggregation layer is to compress high-dimensional feature maps into a D-dimensional feature vector, serving as a global descriptor. Early aggregation layers, including NetVLAD [22], GeM [23] and ConvAP [24], are primarily designed to be integrated with CNN-based backbone networks for enhanced feature representation. With the exceptional performance exhibited by the DINOv2, researchers also introduced novel aggregation techniques, such as SALAD [25], specifically designed for networks utilizing DINOv2 as their backbone. More recently, BoQ [26] introduced an attention-based aggregation method, which leverages learned queries to enhance feature aggregation. However, most current VPR research primarily focuses on outdoor scenarios. Recent studies, such as AnyLoc [18] and VLAD-BuFF [21], highlighted the poor generalization of outdoor-trained models to indoor and structured environments.
@@ II-B Text-based Localization and Place Recognition
In recent years, text-based localization methods have emerged as a promising solution for achieving robust localization in challenging environments. The high-level semantic features of scene text demonstrate remarkable invariance to illumination changes and viewpoint variations. Moreover, these features offer effective discriminative capabilities in highly repetitive scenes. TextPlace [10] is the first to utilize scene text for VPR in the context of topological localization. It demonstrates that scene text can effectively mitigate the challenges posed by extreme visual appearance variations in urban environments. TextSLAM [11] is the first work that tightly integrates text semantic information for loop closure detection (LCD) in visual SLAM. Recently, TextLCD [13] integrated a multi-modal LCD method based on explicit scene text into LiDAR SLAM frameworks and introduced a text-based repetitive scene dataset. However, these methods have two main limitations: (1) They treat scene text spotting as an independent module, leading to a decrease in system computational efficiency. (2) All of these methods rely on a predefined dictionary to filter key text, which significantly limits their adaptability to different environments. Furthermore, most existing methods depend on spatial-temporal constraints to construct text maps, which are not applicable in our setting.
@@ III Methodology
Figure 2: An overview of TextInPlace framework. TextInPlace is (a) a local parameter sharing network with two branches. (b) The Visual Place Recognition Branch extracts global descriptors for coarse global retrieval. (c) The Scene Text Spotting Branch detects and recognizes scene text in query and database images. Finally, re-ranking is performed based on the text similarity between the query image and the retrieved images.
@@ III-B The Visual Place Recognition Branch
Specifically, inspired by the Bag-of-Queries (BoQ) block [26], we introduce a set of learnable queries that interact with the multi-channel feature maps $F_{\text{VPR}}$ through cross-attention. Each block contains a fixed set of learnable queries, denoted as $Q$. The process begins by applying self-attention to $Q$:
@@ III-D Verification based on Discriminative Scene Text
Once the query descriptor $O$ is extracted, a K-nearest neighbor (KNN) search is performed in the existing feature database $D$ based on Euclidean distance. The top-K results are selected. Subsequently, the text $T$ extracted from the query image is compared with the text from the top-K results. It is important to note that certain text elements in scenes, such as ”Fire Hydrant” or ”Emergency Exit”, may appear frequently across various locations, leading to potential perceptual ambiguity. To address this issue, we have two solutions.
@@ III-D1 Rule-based Text Filter
In indoor environments, numerical identifiers, such as door numbers, are commonly used to distinguish repetitive structures. Consequently, our rule-based text filter identifies and filters discriminative text by detecting the presence of numerical digits within the text.
@@ III-D2 LLM-based Text Filter
Due to the vast commonsense knowledge of the LLM, it can handle complex textual information more flexibly. As shown in Fig. 3, when provided with a prompt, the LLM can automatically discern discriminative text. Specifically, our implementation utilizes the cost-effective GPT-4o-min API to achieve this functionality. However, despite its flexibility and generalizability, LLM-based text filtering is less efficient due to the requirement of remote access.
@@
When discriminative text is filtered, the re-ranking and verification are conducted using a string similarity score, defined as follows:
@@ III-E1 Indoor VPR Dataset for Fine-tuning
where $M$ (in meters) and $\alpha$ (in degrees) are two parameters that define the spatial and angular boundaries of each class.
@@ III-E2 Loss function
where $\alpha$, $\beta$, and $\lambda$ are hyperparameters that regulate the weighting of the positive and negative components within the loss function.
@@ IV The Maze-with-Text Dataset
To the best of our knowledge, TextLCD [13] is the most suitable publicly available dataset for evaluating our framework, as it features repetitive structures and scene text. However, the scale of the TextLCD dataset is limited, as it was collected from only three floors. Thus, it fails to capture the perceptual confusion caused by repetitive structures across multiple floors. Moreover, to simplify text spotting, a significant portion of the text images were captured at close range and from a frontal perspective, which does not reflect the environmental structure. This also does not align with the typical scenarios encountered during robot navigation.
To address these limitations and inspired by the semi-automatic annotation method of NYC-Indoor-VPR [27], we develop the Maze-with-Text dataset within a university teaching building. First, we use a handheld Insta360 One X3 spherical camera to record videos with a resolution of $3840\times 1920$ across five floors. On each floor, two videos are captured along approximately the same trajectory at different times, serving as the query and database sets, respectively. These panoramic videos are then processed using OpenVSLAM [33] to generate a set of keyframes and topological trajectories. Finally, the turning points of the topological trajectories are manually matched to align the keyframe poses within the same coordinate frame. To avoid erroneous loop closure detection, we only enable it when the trajectory exhibits significant crossings.
@@
The trajectories of the dataset sequences are illustrated in Fig. 4. Keyframes are used as the images in the dataset. For each equirectangular projection image, we crop it into four perspective images with a resolution of $640\times 480$ with a $90\degree$ field of view. In addition, a text detector is employed to identify and filter images containing scene text, which are then used as query images. Tab. I provides additional details of our dataset.
TABLE I: Number of images of our MAZE-WITH-TEXT dataset
@@ V EXPERIMENTS
This section reports detailed experimental settings and evaluation results on two indoor test datasets with repetitive structures, Maze-with-Text and TextLCD dataset [13].
@@ V-A Implementation details
The proposed method is implemented using PyTorch [34] with model training conducted on two NVIDIA RTX 3090 GPUs, while testing is performed on a single GPU. We employ ResNet-50 [30] backbone for feature extraction, and DPText-DETR [35] is employed for text detection, while DiG [36] is adopted for text recognition. Our method defaults to using a rule-based text filter, while the variant utilizing a LLM-based text filter is designated as TextInPlace-L.
@@
TABLE II: Performance comparison on our Maze-with-Text dataset. The best is highlighted in bold and the second-best is underlined. We report Recall@1 and Recall@5 for each method.
@@ V-B Comparison with State-of-the-art
Table II shows the the comparison of the proposed method with previous SOTA methods, including ten 1-stage models and four 2-stage re-ranking models. Since some text-based methods, like TextPlace [10], not being open-source and most relying on spatial-temporal constraints (e.g., odometry) to build text maps, we do not directly compare with them, but we discuss the representative method TextLCD [13] in the next experiment. For all comparative methods, we adopted the well-trained model provided by their official sources. Due to the constraints on input image resolution, the test image resolution for methods employing the DINOv2 as the backbone and MixVPR is set to $224\times 224$ and $320\times 320$, respectively. All 2-stage methods re-rank the top-100 candidates to improve retrieval performance.
Compared to several recent SOTA approaches, our method demonstrates superior performance. TextInPlace-L exhibits a marginal performance advantage over TextInPlace, underscoring that LLM-based text filter, which dynamically adapt their filtering strategies according to the input text, are more effective at filtering discriminative text than static rule-based text filter. Previous VPR methods often experience significant performance degradation due to severe perceptual confusion when encountering indoor repetitive structures. Notably, methods using the visual foundation model DINOv2 as the backbone network fail to significantly outperform those based on the ResNet-50 backbone network. 2-stage methods that employ local feature-based re-ranking frequently struggle to handle repetitive structures effectively, primarily because of the high similarity among visual elements in the scene. Our method addresses this challenge by leveraging scene text as positional clues, enabling robots to distinguish such scenes in a human-like manner.
To further validate our experimental findings, we conduct comparative evaluations on a public dataset. Specifically, we selected sequences seq1, seq2, and seq4 from the TextLCD dataset [13]. The initial segments of each trajectory serves as the database, while the overlapping subsequent trajectory served as queries. To more closely align with the application of loop closure detection, we select one query image every 3 seconds. It means that scene text may not always be present in the query images.
As shown in Table III, our method achieves the most significant improvement on seq2, surpassing BoQ by 5.7% in Recall@1. On seq1 and seq4, it maintains performance comparable to that of BoQ. Both the 2-stage method, Patch-NV, and the DINOv2-based method, SALAD, demonstrate inferior performance compared to TextInPlace.
TABLE III: Performance comparison on the TextLCD test dataset.
@@
Patch-NV VGG-16 75.7 87.6 69.6 83.7 93.0 93.0
SALAD DINOv2-B 80.4 85.0 56.0 66.7 84.5 90.1
BoQ ResNet-50 84.1 88.8 71.6 77.3 93.0 94.0
@@
Since TextLCD is a multi-modal LiDAR-based LCD method, a direct comparison with our visual approach is not feasible. However, the original TextLCD paper [13] shows that its recall is lower than SALAD, which we attribute to its reliance on text-derived semantic and geometric information, neglecting appearance information. In contrast, our method incorporates global descriptors as prior information and uses discriminative text for spatial verification, outperforming SALAD. Therefore, we argue TextInPlace theoretically outperforms TextLCD in repetitive indoor environments.

### v2

@@ Abstract
Visual Place Recognition (VPR) is a crucial capability for long-term autonomous robots, enabling them to identify previously visited locations using visual information. However, existing methods remain limited in indoor settings due to the highly repetitive structures inherent in such environments. We observe that scene texts frequently appear in indoor spaces and can help distinguish visually similar but different places. This inspires us to propose TextInPlace, a simple yet effective VPR framework that integrates Scene Text Spotting (STS) to mitigate visual perceptual ambiguity in repetitive indoor environments. Specifically, TextInPlace adopts a dual-branch architecture within a local parameter sharing network. The VPR branch employs attention-based aggregation to extract global descriptors for coarse-grained retrieval, while the STS branch utilizes a bridging text spotter to detect and recognize scene texts. Finally, the discriminative texts are filtered to compute text similarity and re-rank the top-K retrieved images. To bridge the gap between current text-based repetitive indoor scene datasets and the typical scenarios encountered in robot navigation, we establish an indoor VPR benchmark dataset, called Maze-with-Text. Extensive experiments on both custom and public datasets demonstrate that TextInPlace achieves superior performance over existing methods that rely solely on appearance information. The dataset, code, and trained models are publicly available at https://github.com/HqiTao/TextInPlace.
@@ I INTRODUCTION
However, state-of-the-art (SOTA) VPR methods often struggle in indoor scenarios, suggesting that indoor VPR remains an underexplored research area. Moreover, indoor VPR presents a more convenient and cost-effective alternative compared to traditional indoor localization methods that rely on WiFi or Radio Frequency Identification (RFID) [8]. This highlights the importance of further research and innovation to develop robust and efficient VPR techniques tailored specifically for indoor environments. We argue that the suboptimal performance of current VPR methods in indoor environments can be primarily attributed to the repeated structure caused by the repetitive and symmetrical designs of man-made buildings. Unlike outdoor environments, indoor spaces often exhibit high visual similarity across different locations (see Fig. 1), particularly in multi-floor buildings where architectural designs, floor layouts, and decorative elements are often highly repetitive [9, 10, 11]. Geometry-based verification approaches are susceptible to false positive correspondences in environments with repetitive structures [12], which significantly undermines the reliability of place recognition in unmanned systems.
@@
Inspired by human cognitive patterns, several recent studies [13, 14, 15, 16, 17] proposed leveraging the high-level semantic information embedded in scene texts for robot localization. However, these methods also typically rely on off-the-shelf Scene Text Spotting (STS) technology, which frequently becomes a computational bottleneck. Furthermore, they mainly use a predefined dictionary of specific words to filter out discriminative texts, and this limits their adaptability to diverse environments and requires significant manual effort. Additionally, existing text-based localization datasets, which are collected with a focus on scene texts, remain limited in number. Most of these datasets deliberately capture scene texts from close distances and frontal viewpoints to facilitate text spotting. However, such data collection strategies do not accurately reflect the typical visual conditions encountered in robot navigation within real-world environments.
@@
To enhance the applicability of VPR in indoor settings, we propose TextInPlace, a unique dual-branch VPR framework that harnesses textual clues for robust spatial verification in highly repetitive indoor environments. Specifically, TextInPlace performs coarse-grained retrieval based on the global descriptor derived from the VPR branch, and executes scene text spotting via a dedicated STS branch applied to both query and retrieved images. Discriminative texts are efficiently filtered out based on predefined rules or, more generally, using a large language model (LLM). It then establishes textual semantic correspondence by matching these discriminative scene texts, enabling explainable re-ranking of the top-K candidate images initially retrieved by VPR branch. Moreover, the VPR branch and STS branch are combined with a local parameter sharing backbone for optimizing computational efficiency. To address the limitations of existing datasets, we use a panoramic camera to establish an indoor VPR benchmark dataset, called Maze-with-Text. The panoramic images are then cropped to generate VPR data. In addition to scene texts, most images in the dataset also capture the structural details of the surrounding environment.
@@
  • We develop an LLM-based text filter that flexibly and generically identifies discriminative texts.
  • We establish Maze-with-Text, an indoor VPR benchmark dataset to highlight the challenge of repetitive structures in multi-floor buildings, with scene texts appearing at various distances and perspectives, thus more closely reflecting robot navigation scenarios.
@@ II-A Visual Place Recognition Network
The VPR network consists mainly of backbone architecture and aggregation layer. Both CNN-based [3] [4] [18] and Transformer-based [19] [20] feature extraction backbones have been widely employed in VPR research, demonstrating their effectiveness. Recently, researchers adopted the vision foundation model DINOv2 [21] as a backbone, exploring both train-free [22] and fine-tuning [23] [24] approaches, and achieving remarkable capabilities. However, vision foundation models are computationally intensive, and relying on large backbones may lead researchers to overlook innovations in other aspects of the VPR problem [25]. The role of the aggregation layer is to compress high-dimensional feature maps into a feature vector, serving as a global descriptor. Early aggregation layers, including NetVLAD [26], GeM [27] and ConvAP [28], are primarily designed to be integrated with CNN-based backbone networks for enhanced feature representation. With the exceptional performance exhibited by the DINOv2, researchers also introduced novel aggregation techniques, such as SALAD [29], specifically designed for networks utilizing DINOv2 as their backbone. More recently, BoQ [30] introduced an attention-based aggregation method, which leverages learned queries to enhance feature aggregation. However, most current VPR research primarily focuses on outdoor scenarios. Recent studies, such as AnyLoc [22] and VLAD-BuFF [25], highlighted the poor generalization of outdoor-trained models to indoor and structured environments.
@@ II-B Text-based Localization and Place Recognition
In recent years, text-based localization methods have emerged as a promising solution for achieving robust localization in challenging environments. The high-level semantic features of scene texts demonstrate remarkable invariance to illumination changes and viewpoint variations. Moreover, these features offer effective discriminative capabilities in highly repetitive scenes. TextPlace [14] is the first to utilize scene texts for VPR in the context of topological localization. It demonstrates that scene text can effectively mitigate the challenges posed by extreme visual appearance variations in urban environments. TextSLAM [15] is the first work that tightly integrates texts semantic information for loop closure detection (LCD) in visual SLAM. Recently, TextLCD [17] integrated a multi-modal LCD method based on explicit scene texts into LiDAR SLAM frameworks and introduced a text-based repetitive scene dataset. However, these methods have two main limitations: (1) They treat scene text spotting as an independent module, leading to a decrease in system computational efficiency, and (2) most of these methods rely on a predefined dictionary to filter key texts, which significantly limits their adaptability to different environments. Furthermore, most existing methods depend on spatial-temporal constraints to construct text maps, whereas the VPR problem typically operates without temporal constraints.
@@ III Methodology
Figure 2: An overview of TextInPlace framework. TextInPlace is (a) a local parameter sharing network with two branches. (b) The Visual Place Recognition Branch extracts global descriptors for coarse global retrieval. (c) The Scene Text Spotting Branch detects and recognizes scene texts in query and database images. Finally, re-ranking is performed based on the text similarity between the query image and the retrieved images.
@@ III-B The Visual Place Recognition Branch
Specifically, we follow the aggregation block described in Bag-of-Queries (BoQ) [30] and introduce a set of learnable queries that interact with the multi-channel feature maps $F_{\text{VPR}}$ through cross-attention. Each block contains a fixed set of learnable queries, denoted as $Q$. The process begins by applying self-attention to $Q$:
@@ III-D Verification based on Discriminative Scene Text
Once the query descriptor $O$ is extracted, a K-nearest neighbor (KNN) search is performed in the existing feature database $D$ based on Euclidean distance. The top-K results are selected. Subsequently, the scene texts $T$ extracted from the query image is compared with the texts from the top-K results. It is important to note that certain text elements in scenes, such as ”Fire Hydrant” or ”Emergency Exit”, may appear frequently across various locations, leading to potential perceptual ambiguity. To address this issue, we propose two solutions.
@@ III-D1 Rule-based Text Filter
In indoor environments, numerical identifiers, such as door numbers, are commonly used to distinguish repetitive structures. Consequently, our rule-based text filter identifies and filters discriminative texts by detecting the presence of numerical digits within the texts.
@@ III-D2 LLM-based Text Filter
Due to the vast commonsense knowledge of the LLM, it can handle complex scene texts more flexibly. As shown in Fig. 3, when provided with a prompt, the LLM can automatically discern discriminative texts. Specifically, our implementation utilizes the GPT-4o-mini API for LLM-based text filter.
@@
When discriminative texts are filtered, the re-ranking and verification are conducted using a string similarity score, defined as follows:
@@ III-E1 Indoor VPR Dataset for Fine-tuning
where $M$ and $\alpha$ are two parameters that define the spatial and angular boundaries of each class.
@@ III-E2 Loss function
where $\alpha$, $\beta$, and $\lambda$ are hyperparameters that regulate the weighting of the positive and negative components.
@@ IV The Maze-with-Text Dataset
To the best of our knowledge, TextLCD [17] is the most suitable publicly available dataset for evaluating our framework, as it features repetitive structures and scene texts. However, the scale of the TextLCD dataset is limited, as it was collected from only three floors. Thus, it fails to capture the perceptual confusion caused by repetitive structures across multiple floors. Moreover, to simplify text spotting, a significant portion of the text images were captured at close distances and frontal viewpoints, which does not reflect the environmental structure. This also does not align with the typical scenarios encountered during robot navigation.
To address these limitations and inspired by the semi-automatic annotation method of NYC-Indoor-VPR [31], we develop the Maze-with-Text dataset within a university teaching building. First, we use a handheld Insta360 One X3 spherical camera to record videos at $3840\times 1920$ resolution across five floors. The higher resolution, relative to that of NYC-Indoor-VPR ($1920\times 960$), facilitates clearer capture of scene texts for improved detection and recognition. On each floor, two videos are captured along approximately the same trajectory at different times, serving as the query and database sets, respectively. These panoramic videos are then processed using OpenVSLAM [37] to generate a set of keyframes and topological trajectories. Finally, the turning points of the topological trajectories are manually matched to align the keyframe poses within the same coordinate frame. To avoid false positive loop closure detection, it is manually enabled only when the trajectory exhibits clear intersections.
@@
The trajectories of the dataset sequences are illustrated in Fig. 4. Keyframes are used as the images in the dataset. Each equirectangular projection image is cropped into four perspective images with a resolution of $640\times 480$ and a $90\degree$ field of view (FOV). In addition, a text detector is employed to identify and retain images containing scene texts, which are then used as query images. Table I provides the number of images in the Maze-with-Text dataset.
TABLE I: Number of images in the Maze-with-Text dataset
@@ V EXPERIMENTS
This section reports detailed experimental settings and evaluation results on two indoor test datasets with repetitive structures, namely Maze-with-Text and TextLCD dataset.
@@ V-A Implementation details
The proposed method is implemented using PyTorch [38] with model training conducted on two NVIDIA RTX 3090 GPUs, while testing is performed on a single GPU. We employ ResNet-50 [34] backbone for feature extraction. DPText-DETR [39] is employed for text detection, while DiG [40] is adopted for text recognition. Our method defaults to using a rule-based text filter, while the variant utilizing a LLM-based text filter is designated as TextInPlace-L.
@@
TABLE II: Performance comparison on the Maze-with-Text dataset. Best, Second Best, and Third Best results are highlighted. We report Recall@1 and Recall@5 for each method.
@@ V-B Comparison with State-of-the-art
Table II shows the the comparison of the proposed method with previous SOTA methods, including nine 1-stage models: NetVLAD [26], SFRS [42], CosPlace [3], EigenPlaces [4], ConvAP [28], MixVPR [41], BoQ [30], SALAD [29], CricaVPR [24], and four 2-stage re-ranking models: Patch-NV [43], TransVPR [19], R2Former [20], SelaVPR [23]. Since some text-based methods, like TextPlace [14], are not open-source and mostly rely on spatial-temporal constraints (e.g., odometry) to build text maps, we do not directly compare them with our method. For all comparative methods, we adopted the well-trained model provided by their official sources. Due to the constraints on input image resolution, the test image resolution for methods employing the DINOv2 as the backbone and MixVPR is set to $224\times 224$ and $320\times 320$, respectively. All 2-stage methods re-rank the top-100 candidates to improve retrieval performance.
Compared to several recent SOTA approaches, our method demonstrates superior performance. TextInPlace-L exhibits a marginal performance advantage over TextInPlace, underscoring that LLM-based text filter, which dynamically adapt their filtering strategies according to the input texts, are more effective at filtering discriminative texts than static rule-based text filter. Previous VPR methods often experience significant performance degradation due to severe perceptual confusion when encountering indoor repetitive structures. Notably, methods using the visual foundation model DINOv2 as the backbone network fail to significantly outperform those based on the ResNet-50 backbone network. 2-stage methods that employ local feature-based re-ranking struggle to handle repetitive structures effectively, primarily because of the high similarity among visual elements in the scene. Our method addresses this challenge by leveraging scene texts as positional clues, enabling robots to distinguish such scenes in a human-like manner.
To further validate our experimental findings, we conduct comparative evaluations on a public dataset. Specifically, we selected sequences seq1, seq2, and seq4 from the TextLCD dataset [17]. The initial segments of each trajectory serves as the database, while the overlapping subsequent trajectory served as queries. To more closely align with the application of loop closure detection, we select one query image every 3 seconds. It means that scene texts may not always be present in the query images.
As shown in Table III, our method achieves the significant improvement on seq2, surpassing BoQ by 5.7% in Recall@1. On seq1 and seq4, it maintains performance comparable to that of BoQ. Both the 2-stage method, Patch-NV, and the DINOv2-based method, SALAD, demonstrate inferior performance compared to TextInPlace.
TABLE III: Performance comparison on TextLCD test dataset. The best is highlighted in bold and the second-best is underlined.
@@
Patch-NV [43] VGG-16 75.7 87.6 69.6 83.7 93.0 93.0
SALAD [29] DINOv2-B 80.4 85.0 56.0 66.7 84.5 90.1
BoQ [30] ResNet-50 84.1 88.8 71.6 77.3 93.0 94.0
@@
Since TextLCD is a multi-modal LiDAR-based LCD method, a direct comparison with our visual approach is not feasible. However, the original TextLCD paper [17] shows that its recall is lower than SALAD, which we attribute to its reliance on text-derived semantic and geometric information, neglecting appearance information. In contrast, our method incorporates global descriptors as prior information and uses discriminative texts for spatial verification, outperforming SALAD. Therefore, we argue TextInPlace theoretically outperforms TextLCD in repetitive indoor environments.